# Graph Sampling for Scalable and Expressive Graph Neural Networks on Homophilic Graphs

### v3

@@ Abstract
Index Terms:
graph signal processing, graph sampling, graph neural networks, transferability, homophily.
@@ I Introduction
GNNs are deep neural networks tailored to network data which have shown great empirical performance in several graph machine learning tasks [1, 2, 3, 4]. This is especially true in graph signal processing problems—such as recommender systems on product similarity networks [5], or attribution of research papers to scientific domains [6]—in which GNNs’ invariance and stability properties [7, 8] play a key role.
@@
Yet, in practice most successful applications of GNNs are limited to graphs of moderate size. The sheer size of many modern networks, typically in the order of several millions, frequently makes these models impractical to train. Good results have been seen by leveraging the GNN’s transferability property [9, 10], which states that a GNN with fixed weights produces similar outputs on large enough graphs belonging to the same “family”, e.g., the same random graph model. This property allows training the GNN on a graph of moderate size, and transferring it for inference on the large graph.
@@
The transferability of GNNs is closely related to their convolutional parametrization, and is a consequence of graph convolutions converging on sequences of graphs that converge to a common graph limit [11]. Under certain assumptions on the type of limit, and on how the graphs converge to (or are sampled from) them, it is possible to obtain non-asymptotic error bounds inversely proportional to the sizes of the graphs. Such bounds are then used to inform practical considerations, such as the minimum graph size on which to train a GNN to meet a maximum transference error. Once this is determined, the training graph is obtained by sampling a subgraph of the appropriate size at random from the large graph.
@@
Learning GNNs on randomly subsampled graphs works reasonably well on average but, for models trained on small samples, there is large variance in performance, and the worst-case performance can be quite low; see Figure 3. While this is a natural consequence of subsampling any type of data, on graphs issues are exacerbated by the fact that random subgraphs often have disconnected components and isolated nodes. This leads to rank reduction in graph matrix representations, which in turn affects the expressive power of GNNs [12]. Indeed, the loss of rank is explicit for the graph Laplacian, in which the multiplicity of the zero eigenvalue indicates the number of connected components in the graph.
@@
Sampling algorithms better at preserving matrix rank—such as spectral algorithms, e.g. [13, 14]—or connectivity—such as local algorithms, e.g., breadth-first search[15]—exist, however, spectral methods have high computational complexity, and local methods focus too much on specific regions of the graph and fail to capture global structure.
@@
In this paper, we identify a property that allows graphs to be sampled efficiently without restricting to local regions: feature homophily. Specifically, let $G=(V,E)$ be a graph with node features $X\in{\mathbb{R}}^{|V|\times d}$. This graph is feature-homophilic if, given that $(i,j)$ is an edge in $G$, the normalized features $\hat{X}[i,:]$ and $\hat{X}[j,:]$ are close. Our first contribution is to introduce a novel definition of feature homophily based on the graph Laplacian. Then, we show that, by sampling nodes to minimize the trace of the correlation matrix $XX^{T}$, it is possible to improve the trace of the graph Laplacian, which provides a measure of connectivity of the subsampled graph.
@@ II Preliminaries
The graph Laplacian plays an important role in graph signal processing (GSP) [16, 17], as it allows defining the notion of total variation of a signal $x$. Explicitly, the total variation of $x$ is defined as $TV(x)=x^{T}{\mathbf{L}}x$ [18]. Let ${\mathbf{L}}={\mathbf{V}}\boldsymbol{\Lambda}{\mathbf{V}}^{T}$ be the Laplacian eigendecomposition, where $\boldsymbol{\Lambda}$ is a diagonal matrix with eigenvalues ordered as $\lambda_{1}\leq\ldots\leq\lambda_{n}$ and ${\mathbf{V}}$ is the corresponding eigenvector matrix. For unit-norm signals, it is easy to see that the maximum total variation is $\lambda_{n}$, the largest Laplacian eigenvalue, and the minimum total variation is $\lambda_{1}=0$, which corresponds to the all-ones eigenvector. Therefore, the Laplacian eigenvalues can be interpreted as graph frequencies, and the eigenvectors as these frequencies’ respective oscillation modes.
@@ II-A Graph Neural Networks
GNNs are deep learning models specifically designed for graph-structured data, where each layer consists of two components: a bank of convolutional filters and a nonlinear activation function. A graph convolutional filter is the extension of a standard convolutional filter to graph data. More specifically, it consists of a shift-and-sum operation of a signal $x$ on the graph $G$, which is captured by a matrix ${\mathbf{S}}\in{\mathbb{R}}^{n\times n}$ encoding the sparsity pattern of the graph, i.e., ${\mathbf{S}}[i,j]\neq 0$ if and only if $(i,j)\in E$ or $i=j$; typical choices are ${\mathbf{S}}={\mathbf{A}}$ or ${\mathbf{S}}={\mathbf{L}}$ [19]. The graph shift operator operates on $x$ as ${\mathbf{S}}x$.
@@
Given any choice of ${\mathbf{S}}$, the graph convolution is defined as $y=\sum_{k=0}^{K-1}h_{k}{\mathbf{S}}^{k}x$, where $h_{0},\ldots,h_{K-1}$ are the filter coefficients or taps. More generally, for $X\in{\mathbb{R}}^{n\times d}$ and $Y\in{\mathbb{R}}^{n\times f}$ we can define the convolutional filterbank [4]
@@
where ${\mathbf{H}}_{k}\in{\mathbb{R}}^{d\times f}$, $10\leq k\leq K-$ 1.
The $\ell$th layer of a GNN is given by [4]
@@
with $\sigma:{\mathbb{R}}\to{\mathbb{R}}$ an entry-wise nonlinearity (e.g., the ReLU or sigmoid). At layer $\ell=1$, $X_{0}$ is the input data, and the last layer output $X_{L}$ is the output $Y$ of the GNN. For succinctness, in the following we will represent the whole $L$-layer GNN as a map $Y=\Phi(X,G;{\mathcal{H}})$ with =${\mathcal{H}}=\{{\mathbf{H}}_{\ell k}\}_{\ell,k}$.
@@ II-B Transferability of GNNs
A graphon is a bounded, symmetric, measurable function ${\mathbf{W}}:[0,1]^{2}\to[0,1]$ [20, 21]. Graphs can be sampled from ${\mathbf{W}}$ by sampling nodes $u_{1},\ldots,u_{n}$ from $[0,1]$, and sampling edges $(u_{i},u_{j})$ with probability ${\mathbf{W}}(u_{i},u_{j})$. The graph limit interpretation is more nuanced but has to do with the fact that, on sequences of graphs converging to ${\mathbf{W}}$, the densities of certain “motifs”, e.g., triangles, also converge. For graphs associated with the same graphon, we have the following transferability theorem.
Theorem II.1 (GNN transferability, simplified [22]).
Let $\Phi$ be a GNN with fixed coefficients, and $G_{n}$, $G_{m}$ graphs with $n$ and $m$ nodes sampled from a graphon $\mathbf{W}$. Under mild conditions, $\|{\Phi}(G_{n})-{\Phi}(G_{m})\|=\mathcal{O}(n^{-1}+m^{-1})$ w.h.p.
@@
Figure 1: Adjusted Laplacian trace versus graph subsampling rate. The adjusted trace is the subsampled graph Laplacian trace normalized by the number of sampled nodes. Boxplots indicate the mean and standard error of the trace obtained from 50 rounds of random node subsampling; red dots are the trace of subgraphs generated using our sampling heuristic (Algorithm 1). Trace preservation is almost always better for our algorithm compared with the average of random subsampling, except for Cora at 75% sample rate.
@@ II-C Expressive Power of GNNs
While GNNs achieve remarkable performance in many graph machine learning tasks, they have fundamental limitations associated with their expressive power [23, 24]. In GSP problems specifically, the expressivity of a GNN is constrained by the expressivity of the graph convolution, which in turn is constrained by the rank of the graph shift operator [12]. This is demonstrated in the following proposition.
@@
Figure 3: Test accuracy achieved by GNN on full graph versus training graph subsampling rate. Boxplots indicate the spread of the accuracy realized by GNNs trained on 50 random node-induced subgraphs; red dots are the test accuracy of GNNs trained on subgraphs produced by our heuristic (Algorithm 1). The feature-based heuristic yields better results across the board, except for CiteSeer at 25% sample rate.
@@ III Feature homophily and homophily-based sampling
We start by introducing the notion of feature homophily. In order to make this definition compatible for graphs of different sizes and with different features features, we first need to normalize $X\in\mathbb{R}^{n\times d}$ along both the feature and node dimensions. Explicitly, let $\vec{\mu}=(\mu_{1},\cdots,\mu_{d})^{T}\in\mathbb{R}^{d}$ be the mean feature vector and $\vec{\sigma}=(\frac{1}{\sigma_{1}},\cdots,\frac{1}{\sigma_{d}})^{T}\in\mathbb{R}^{d}$ the standard deviation vector. We define the normalized graph feature matrix as:
$\hat{X}=(X-\mathbf{1}_{n}\vec{\mu}^{T})\odot(\frac{1}{\sqrt{d}}\cdot\mathbf{1}_{n}\vec{\sigma}^{T})\text{.}$ (3)
I.e., $\hat{X}[i,j]=\frac{X[i,j]-\mu_{j}}{\sqrt{d}\sigma_{j}}$.
@@
Since ${\mathbf{L}}$ is positive semidefinite and the trace of the outer product is the product of traces, it is ready to see, by Cauchy-Schwarz, that $h_{G}\leq 0$ for any undirected graph $G$. The larger the feature homophily $h_{G}$, i.e., the closer it is to 00, the higher the alignment of the data $X$ (or, more precisely, of its principal components) with the low-frequency eigenvectors of ${\mathbf{L}}$—which account for most of the graph’s global structure such as its connected components and clusters/communities. Thus, the data is informative with regards to the graph. On the other hand, highly negative values of $h_{G}$ indicate strong alignment of $\hat{X}$ with high-frequency eigenvectors, which tend to be noisier and less descriptive of the graph structure.
@@
$\text{tr}({\mathbf{L}})\geq-\frac{nh_{G}}{\text{tr}(\hat{X}\hat{X}^{T})}$ (5)
@@ III-A Sampling Heuristic for Homophilic Graphs
Although simple, the result in Proposition III.2 has important implications for feature-homophilic graph sampling. Assume we start removing nodes from $G$ according to the diagonal entries of $XX^{T}$ sorted in decreasing order, so that the denominator on the right-hand side of (5) becomes progressively smaller. We prioritize removing nodes corresponding to the largest diagonal elements because, in order to tighten the lower bound of ${tr}({\mathbf{L}})$, we should enforce the denominator ${tr}(\hat{X}\hat{X}^{T})^{2}$ to be as small as possible, given that the graph is homophilic. Hence, deleting from largest to smallest makes sense. This idea is formalized in Algorithm 1.
@@
On sparse graphs ($m\ll n^{2}$) with moderate feature dimension $d$, Algorithm 1 is cheaper than direct maximization of $tr({\mathbf{L}})$, which requires $O((1-\gamma)n^{2})$ computations—$O(n)$ node degree computations $(1-\gamma)n$ times, as node degrees change each time a node is removed from $G$. In fact, this is an underestimation, as it does not factor in the cost of breaking ties across nodes with the same degree. Algorithm 1 is also notably cheaper than spectral algorithms such as [13, 14] which are inspired by E-optimal sampling and, without exhaustive search, require greedy routines with complexity at least $O(\gamma nm)$. Another advantage of Algorithm 1 is that it does not require sequential execution, unlike maximization of $tr({\mathbf{L}})$ and [13, 14], which cannot be parallelized.
@@ III-B Connections with Orthogonality, Leverage Scores Sampling, and Further Discussion
The feature homophily score can also be rewritten as a weighted $\ell_{1}$-norm of the adjacency matrix, where the weight of each edge $A_{ij}$ is the squared Euclidean distance between feature vectors $X_{i}$ and $X_{j}$ [25]. This formulation connects homophily with graph sparsity, suggesting that our sampling heuristic could serve as a structured sparsification technique for large-scale graphs similar to graph sparsification techniques based on effective resistances [26].
The sampling method we propose in Algorithm 1 has interesting connections with leverage scores sampling approaches used in regression models. In classical settings, leverage scores quantify the influence of data points on the regression outcome. However, unlike traditional leverage score sampling, which retains influential, high-leverage points, our approach removes the nodes with the largest scores.
@@
This criterion aligns with the goal of preserving Laplacian rank: since the multiplicity of the zero eigenvalue counts the number of connected components, removing nodes that form separate components or have low connectivity minimally affects the overall structure. Specifically, if a graph initially has $n$ nodes and $k$ connected components (so that $\text{rank}(\mathbf{L})=n-k$), removing an isolated node decreases both $n$ and $k$ by one, preserving the rank at $n-k$. This suggests that our method implicitly filters out nodes that contribute little to the global structure while retaining the essential connectivity properties of the original graph.

### v5

@@ I Introduction
GNNs are deep neural networks tailored to network data which have shown great empirical performance in several graph machine learning tasks [gori2005new, kipf17-classifgcnn, defferrard17-cnngraphs, gama18-gnnarchit]. This is especially true in graph signal processing problems—such as recommender systems on product similarity networks [ruiz2020gnns], or attribution of research papers to scientific domains [hamilton2017inductive]—in which GNNs’ invariance and stability properties [ruiz19-inv, gama19-stability] play a key role.
@@
Yet, in practice most successful applications of GNNs are limited to graphs of moderate size. The sheer size of many modern networks, typically in the order of several millions, frequently makes these models impractical to train. Good results have been seen by leveraging the GNN’s transferability property [ruiz20-transf, levie2019transferability], which states that a GNN with fixed weights produces similar outputs on large enough graphs belonging to the same “family”, e.g., the same random graph model. This property allows training the GNN on a graph of moderate size, and transferring it for inference on the large graph.
@@
The transferability of GNNs is closely related to their convolutional parametrization, and is a consequence of graph convolutions converging on sequences of graphs that converge to a common graph limit [ruiz2020graphonsp]. Under certain assumptions on the type of limit, and on how the graphs converge to (or are sampled from) them, it is possible to obtain non-asymptotic error bounds inversely proportional to the sizes of the graphs. Such bounds are then used to inform practical considerations, such as the minimum graph size on which to train a GNN to meet a maximum transference error. Once this is determined, the training graph is obtained by sampling a subgraph of the appropriate size at random from the large graph.
@@
Learning GNNs on randomly subsampled graphs works reasonably well on average but, for models trained on small samples, there is large variance in performance, and the worst-case performance can be quite low; see Figure 3. While this is a natural consequence of subsampling any type of data, on graphs issues are exacerbated by the fact that random subgraphs often have disconnected components and isolated nodes. This leads to rank reduction in graph matrix representations, which in turn affects the expressive power of GNNs [ruiz2024spectral]. Indeed, the loss of rank is explicit for the graph Laplacian, in which the multiplicity of the zero eigenvalue indicates the number of connected components in the graph.
@@
Sampling algorithms better at preserving matrix rank—such as spectral algorithms, e.g. [chen2015discrete, anis2016efficient]—or connectivity—such as local algorithms, e.g., breadth-first search[alimohammadi2023localgraphlimitsperspective]—exist, however, spectral methods have high computational complexity, and local methods focus too much on specific regions of the graph and fail to capture global structure.
@@
In this paper, we identify a property that allows graphs to be sampled efficiently without restricting to local regions: feature homophily. Specifically, let $G=(V,E)$ be a graph with node features $X\in{\mathbb{R}}^{|V|\times d}$. This graph is feature-homophilic if, given that $(i,j)$ is an edge in $G$, the features X[i,:]{X}[i,:] and X[j,:]{X}[j,:] are close. Our first contribution is to introduce a novel definition of feature homophily based on the graph Laplacian. Then, we show that, by sampling nodes to minimize the trace of the correlation matrix $XX^{T}$, it is possible to improve the trace of the graph Laplacian, which provides a measure of connectivity of the subsampled graph.
@@ II Preliminaries
The graph Laplacian plays an important role in graph signal processing (GSP) [shuman13-mag, sandryhaila13-dspg], as it allows defining the notion of total variation of a signal $x$. Explicitly, the total variation of $x$ is defined as $TV(x)=x^{T}{\mathbf{L}}x$ [ortega2018graph]. Let ${\mathbf{L}}={\mathbf{V}}\boldsymbol{\Lambda}{\mathbf{V}}^{T}$ be the Laplacian eigendecomposition, where $\boldsymbol{\Lambda}$ is a diagonal matrix with eigenvalues ordered as $\lambda_{1}\leq\ldots\leq\lambda_{n}$ and $\mathbf{V}$} is the corresponding eigenvector matrix. For unit-norm signals, it is easy to see that the maximum total variation is $\lambda_{n}$, the largest Laplacian eigenvalue, and the minimum total variation is $\lambda_{1}=0$, which corresponds to the all-ones eigenvector. Therefore, the Laplacian eigenvalues can be interpreted as graph frequencies, and the eigenvectors as these frequencies’ respective oscillation modes.
@@ II-A Graph Neural Networks
GNNs are deep learning models specifically designed for graph-structured data, where each layer consists of two components: a bank of convolutional filters and a nonlinear activation function. A graph convolutional filter is the extension of a standard convolutional filter to graph data. More specifically, it consists of a shift-and-sum operation of a signal $x$ on the graph $G$, which is captured by a matrix ${\mathbf{S}}\in{\mathbb{R}}^{n\times n}$ encoding the sparsity pattern of the graph, i.e., ${\mathbf{S}}[i,j]\neq 0$ if and only if $(i,j)\in E$ or $i=j$; typical choices are ${\mathbf{S}}={\mathbf{A}}$ or ${\mathbf{S}}={\mathbf{L}}$ [segarra17-linear]. The graph shift operator operates on $x$ as ${\mathbf{S}}x$.
@@
Given any choice of $\mathbf{S}$}, the graph convolution is defined as $y=\sum_{k=0}^{K-1}h_{k}{\mathbf{S}}^{k}x$, where $h_{0},\ldots,h_{K-1}$ are the filter coefficients or taps. More generally, for $X\in{\mathbb{R}}^{n\times d}$ and $Y\in{\mathbb{R}}^{n\times f}$ we can define the convolutional filterbank [gama18-gnnarchit]
@@
where ${\mathbf{H}}_{k}\in{\mathbb{R}}^{d\times f}$, $0\leq k\leq K-1$.
[…]
@@
with $\sigma:{\mathbb{R}}\to{\mathbb{R}}$ an entry-wise nonlinearity (e.g., the ReLU or sigmoid). At layer $\ell=1$, $X_{0}$ is the input data, and the last layer output $X_{L}$ is the output $Y$ of the GNN. For succinctness, in the following we will represent the whole $L$-layer GNN as a map $Y=\Phi(X,G;{\mathcal{H}})$ with ${\mathcal{H}}=\{{\mathbf{H}}_{\ell k}\}_{\ell,k}$.
@@ II-B Transferability of GNNs
A graphon is a bounded, symmetric, measurable function ${\mathbf{W}}:[0,1]^{2}\to[0,1]$ [borgs2008convergent, lovasz2012large]. Graphs can be sampled from $\mathbf{W}$} by sampling nodes $u_{1},\ldots,u_{n}$ from $[0,1]$, and sampling edges $(u_{i},u_{j})$ with probability ${\mathbf{W}}(u_{i},u_{j})$. The graph limit interpretation is more nuanced but has to do with the fact that, on sequences of graphs converging to $\mathbf{W}$}, the densities of certain “motifs”, e.g., triangles, also converge. For graphs associated with the same graphon, we have the following transferability theorem.
Theorem II.1 (GNN transferability, simplified [ruiz2021transferability]).
Let $\Phi$ be a GNN with fixed coefficients, and $G_{n}$, $G_{m}$ graphs with $n$ and $m$ nodes sampled from a graphon $\mathbf{W}$. Under mild conditions, w.h.p.,
$\|{\Phi}(G_{n})-{\Phi}(G_{m})\|=\smash{\mathcal{O}\bigg(\sqrt{\frac{\log{n}}{n}}+\sqrt{\frac{\log{m}}{m}}\bigg)}.$
@@
Figure 1: Adjusted Laplacian trace versus graph subsampling rate. The adjusted trace is the subsampled graph Laplacian trace normalized by the number of sampled nodes. Boxplots indicate the median, first and third quartiles, minimum and maximum, and outliers of the trace, obtained from 50 rounds of random node subsampling; red dots are the trace of subgraphs generated using our sampling heuristic (Algorithm 1).
@@ II-C Expressive Power of GNNs
While GNNs achieve remarkable performance in many graph machine learning tasks, they have fundamental limitations associated with their expressive power [xu2018how, chen2019equivalence]. In GSP problems specifically, the expressivity of a GNN is constrained by the expressivity of the graph convolution, which in turn is constrained by the rank of the graph shift operator [ruiz2024spectral]. This is demonstrated in the following proposition.
@@
Figure 3: Test accuracy achieved by GNN on full graph versus training graph subsampling rate. Error bars indicate the standard deviation of the accuracy realized by GNNs trained on 50 random node-induced subgraphs; red dots are the test accuracy of GNNs trained on subgraphs produced by our heuristic (Algorithm 1).
@@ III Feature homophily and homophily-based sampling
We start by introducing the notion of feature homophily. In order to make this definition compatible for graphs of different sizes and with different features, we first need to normalize $X\in\mathbb{R}^{n\times d}$ as
$\hat{X}=\frac{X}{\max_{i,j}|X_{ij}|\sqrt{d}}$ (3)
to ensure all rows have at most unit norm.
@@
Since $\mathbf{L}$} is positive semidefinite and the trace of the outer product is the product of traces, it is ready to see, by Cauchy-Schwarz, that $h_{G}\leq 0$ for any undirected graph $G$. The larger the feature homophily $h_{G}$, i.e., the closer it is to 0, the higher the alignment of the data $X$ (or, more precisely, of its principal components) with the low-frequency eigenvectors of $\mathbf{L}$}—which account for most of the graph’s global structure such as its connected components and clusters/communities. Thus, the data is informative with regards to the graph. On the other hand, highly negative values of $h_{G}$ indicate strong alignment of $\hat{X}$ with high-frequency eigenvectors, which tend to be noisier and less descriptive of the graph structure.
@@
${tr}({\mathbf{L}})\geq-\frac{nh_{G}}{{tr}(\hat{X}\hat{X}^{T})}$ (5)
@@ III-A Sampling Heuristic for Homophilic Graphs
Although simple, the result in Proposition III.2 has important implications for feature-homophilic graph sampling. Assume we start removing nodes from $G$ according to the diagonal entries of $XX^{T}$ sorted in decreasing order, so that the denominator on the right-hand side of (5) becomes progressively smaller. We prioritize removing nodes corresponding to the largest diagonal elements because, in order to tighten the lower bound of ${tr}({\mathbf{L}})$, we should enforce the denominator ${tr}({X}{X}^{T})^{2}$ to be as small as possible, given that the graph is homophilic. Hence, deleting from largest to smallest makes sense. This idea is formalized in Algorithm 1.
@@
On sparse graphs ($m\ll n^{2}$) with moderate feature dimension $d$, Algorithm 1 is cheaper than direct maximization of $tr({\mathbf{L}})$, which requires $O((1-\gamma)n^{2})$ computations—$O(n)$ node degree computations $(1-\gamma)n$ times, as node degrees change each time a node is removed from $G$. In fact, this is an underestimation, as it does not factor in the cost of breaking ties across nodes with the same degree. Algorithm 1 is also notably cheaper than spectral algorithms such as [chen2015discrete, anis2016efficient] which are inspired by E-optimal sampling and, without exhaustive search, require greedy routines with complexity at least $O(\gamma nm)$. Another advantage of Algorithm 1 is that it does not require sequential execution, unlike maximization of $tr({\mathbf{L}})$ and [chen2015discrete, anis2016efficient], which cannot be parallelized.
@@ III-B Connections with Orthogonality, Leverage Scores Sampling, and Further Discussion
The feature homophily score can also be rewritten as a weighted $\ell_{1}$-norm of the adjacency matrix, where the weight of each edge $A_{ij}$ is the squared Euclidean distance between feature vectors $X_{i}$ and $X_{j}$ [kalofolias2016learn]. This formulation connects homophily with graph sparsity, suggesting that our sampling heuristic could serve as a structured sparsification technique for large-scale graphs similar to graph sparsification techniques based on effective resistances [spielman2011graph].
The sampling method we propose in Algorithm 1 has interesting connections with leverage scores sampling approaches used in regression models. In classical settings, leverage scores quantify the influence of data points on the regression outcome. However, unlike traditional leverage score sampling, which retains influential, high-leverage points, our approach removes the nodes with the largest diagonal scores.
@@
This criterion aligns with the goal of preserving Laplacian rank: since the multiplicity of the zero eigenvalue counts the number of connected components, removing nodes that form separate components or have low connectivity minimally affects the overall structure. Specifically, if a graph initially has $n$ nodes and $k$ connected components (so that $\text{rank}(\mathbf{L})=n-k$), removing an isolated node decreases both $n$ and $k$ by one, preserving the rank at nkn-k. This suggests that our method implicitly filters out nodes that contribute little to the global structure while retaining the essential connectivity properties of the original graph.